# Uni-LoRA: One Vector is All You Need

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) has become the de facto parameter-efficient fine-tuning (PEFT) method for large language models (LLMs) by constraining weight updates to low-rank matrices. Recent works such as Tied-LoRA, VeRA, and VB-LoRA push efficiency further by introducing additional constraints to reduce the trainable parameter space. In this paper, we show that the parameter space reduction strategies employed by these LoRA variants can be formulated within a unified framework, Uni-LoRA, where the LoRA parameter space, flattened as a high-dimensional vector space $\mathbb{R}^{D}$, can be reconstructed through a projection from a subspace $\mathbb{R}^{d}$, with $d\ll D$. We demonstrate that the fundamental difference among various LoRA methods lies in the choice of the projection matrix, $P\in\mathbb{R}^{D\times d}$. Most existing LoRA variants rely on layer-wise or structure-specific projections that limit cross-layer parameter sharing, thereby compromising parameter efficiency. In light of this, we introduce an efficient and theoretically grounded projection matrix that is isometric, enabling global parameter sharing and reducing computation overhead. Furthermore, under the unified view of Uni-LoRA, this design requires only a single trainable vector to reconstruct LoRA parameters for the entire LLM – making Uni-LoRA both a unified framework and a “one-vector-only” solution. Extensive experiments on GLUE, mathematical reasoning, and instruction tuning benchmarks demonstrate that Uni-LoRA achieves state-of-the-art parameter efficiency while outperforming or matching prior approaches in predictive performance.
@@ 1 Introduction
To address the aforementioned limitations, we propose Uni-LoRA, a unified framework of LoRA that treats the LoRA parameter space as a high-dimensional vector space and performs a global projection into a shared low-dimensional subspace. We further introduce an efficient and theoretically grounded projection matrix $P\in\mathbb{R}^{D\times d}$, in which each row is a one-hot vector with the index of "1" sampled uniformly from $d$ slots, followed by a column-wise normalization. Conceptually, this corresponds to randomly partitioning all the $D$ parameters of LoRA into $d$ groups and enforce the parameters within each group to share the same value during the training process. We prove that the resulting projection matrix is isometric (or distance-preserving). Empirically, when fine-tuning the GEMMA-7B model, our Uni-LoRA achieves performance comparable to LoRA while training only 0.52M parameters – only 0.0061% of the base model size and 0.26% of the LoRA parameter size. Across multiple benchmarks, our method achieves extreme parameter efficiency, while outperforming or matching the state-of-the-art LoRA variants. Moreover, we show that our projection matrix attains the performance of the Fastfood projection (10.5555/3042817.3042964, ). While Fastfood is a widely used structured projection method with a time complexity of $\mathcal{O}(D\log d)$, our method achieves a significantly lower time complexity of $\mathcal{O}(D)$. Our contributions are summarized as follows:
@@ 3.1 Preliminaries: LoRA and its Parameter-efficient Variants
Tied-LoRA renduchintala-etal-2024-tied and VeRA (kopiczko2024vera, ) augment the LoRA architecture with two extra diagonal matrices. Both methods constrain the weight increment as $\Delta W=\Lambda_{b}P_{B}\Lambda_{d}P_{A}$, where $P_{B}\in\mathbb{R}^{m\times r}$ and $P_{A}\in\mathbb{R}^{r\times n}$ are shared/tied across all LoRA modules, and $\Lambda_{b}\in\mathbb{R}^{m\times m}$ and $\Lambda_{d}\in\mathbb{R}^{r\times r}$ (per LoRA module) are trainable diagonal matrices that can selectively enable or disable columns and rows of $P_{B}$ and $P_{A}$ by scaling, allowing effective, fine-grained adaptation with a minimal number of trainable parameters. The main difference between Tied-LoRA and VeRA is that while $P_{B}$ and $P_{A}$ are trainable in Tied-LoRA, they are randomly initialized and frozen in VeRA, leading to higher parameter efficiency.
@@ 3.3 Analysis of the Projection Matrix
Let $P\in\mathbb{R}^{D\times d}$ be a projection matrix where each row selects exactly one entry uniformly at random to be nonzero, and sets all other entries to zero. Let $n_{j}$ denote the number of nonzero entries in column $j$, and $n_{j}>0$. For column-wise normalization, each nonzero entry in column $j$ is set to $1/\sqrt{n_{j}}$. As a projection matrix, $P$ is isometric. Formally, $\|P(x-y)\|=\|x-y\|,\forall x,y\in\mathbb{R}^{d}$.
@@
Given the construction of $P$ in the theorem statement, we begin by showing that $P^{\top}P=I_{d}$. Consider the $(j,k)$-th entry of $P^{\top}P$: $[P^{\top}P]_{j,k}=\sum_{i=1}^{D}P_{i,j}P_{i,k}.$ Case 1: $j\neq k$. Since each row contains only one nonzero entry, there exists no row $i$ such that both $P_{i,j}$ and $P_{i,k}$ are nonzero. Hence, every term in the summation is zero, and we have $[P^{\top}P]_{j,k}=0$. Case 2: $j=k$. Column $j$ contains $n_{j}$ nonzero entries, each of value $1/\sqrt{n_{j}}$. Thus, $[P^{\top}P]_{j,j}=\sum_{i=1}^{D}P_{i,j}^{2}=n_{j}\cdot\left({1}/{\sqrt{n_{j}}}\right)^{2}=1$. From the analysis above, we have $[P^{\top}P]_{j,k}=1$, if $j=k$ and 0 otherwise, thus $P^{\top}P=I_{d}$. Therefore,
@@ 4 Experiments
In this section, we evaluate our method through a series of experiments. We begin by comparing Uni-LoRA to the state-of-the-art PEFT methods: LoRA, Tied-LoRA, VeRA, LoRA-XS, and FourierFT on the GLUE benchmark. Next, we extend our analysis to mathematical reasoning tasks on Mistral and Gemma models, as well as instruction tuning tasks on Llama2. All our experiments are conducted on a server equipped with 8 NVIDIA A6000 GPUs. For reproducibility, we provide detailed hyperparameters and specifications of computing resources for each experiment in Appendix A.2.
@@ 4.2 Mathematical Reasoning
To evaluate mathematical reasoning capabilities, we fine-tune the Mistral-7B-v0.1 (jiang2023mistral7b, ) and Gemma-7B (gemmateam2024gemmaopenmodelsbased, ) models on the MetaMathQA (yu2023metamath, ) dataset and test them on GSM8K (cobbe2021gsm8k, ) and MATH (hendrycksmath2021, ). We compare our results with the state-of-the-art methods VB-LoRA (li2024vblora, ) and LoRA-XS (bałazy2024loraxslowrankadaptationextremely, ), following their experimental configuration. Table 4 shows that Uni-LoRA demonstrates strong and consistent performance across both Mistral-7B and Gemma-7B backbones. On the MATH benchmark, it achieves scores of 18.18 with Mistral-7B and 28.94 with Gemma-7B, outperforming both LoRA-XS and VB-LoRA baselines. Compared to LoRA-XS, Uni-LoRA yields higher accuracy on both GSM8K and MATH while using less than 70% of the trainable parameters on both backbone models. Compared to VB-LoRA, Uni-LoRA achieves a competitive performance with substantially fewer trainable parameters. Even when comparing in terms of storage cost, Uni-LoRA remains more compact than VB-LoRA because Uni-LoRA only needs to store the subspace vector $\theta_{d}$ and a random seed. Notably, Uni-LoRA uses only 0.52M trainable parameters – only 0.31% of the LoRA parameters on Mistral-7B and 0.26% of the LoRA parameters on Gemma-7B – while achieving the competitive performance of full-parameterized LoRA.
Table 3: Results with the Mistral-7B and Gemma-7B models on the GSM8K and MATH Benchmarks. The best results in each group are shown in bold. VB-LoRA and LoRA-XS results are either reproduced using their reported configurations or directly sourced from their original papers.
@@ 4.4 Ablation Study
Fastfood (10.5555/3042817.3042964, ) is a classical projection method known for its low computational and memory cost and distance-preserving properties. To evaluate the effectiveness of our projection method, we compare it against Uni-LoRA (Fastfood) on four GLUE tasks in terms of both predictive performance and runtime efficiency. Table 6 shows that Uni-LoRA with random projection consistently outperforms Uni-LoRA (Fastfood) in terms of predictive performance across four GLUE tasks, while being significantly faster. For example, on MRPC, it achieves a slightly higher score (91.3 vs. 90.7) with a 65% reduction in training time (9 vs. 26 mins). This efficiency stems from the linear time complexity $\mathcal{O}(D)$ of our projection, compared to Fastfood’s $\mathcal{O}(D\log d)$.
@@
To investigate the impact of global vs. layer-wise (local) projections to our method, we construct a controlled variant as follows. On the RoBERTa-large model, we project the parameters of each layer into its own trainable subspace, with each local projection matrix constructed in the same way as that of Uni-LoRA. To ensure a fair comparison, we set the per-layer subspace dimensionality such that the total number of trainable space dimensionality matches that of Uni-LoRA’s. Table 6 shows that local projection performs worse than global projection on MRPC, SST-2, and QNLI, while both methods achieve comparable performance on CoLA.
@@
To investigate the impact of uniform vs. non-uniform projections to our method, we construct a controlled variant in which two separate projection matrices are used: one projects all LoRA $A$ matrices into the first two-thirds of the subspace dimensions, and the other projects all $B$ matrices into the remaining one-third. Both projection matrices are constructed using the same method as that of Uni-LoRA. Table 6 shows that the uniform projection consistently outperforms the non-uniform projection across all the tasks.
@@ 5 Conclusion
Limitations and broader impacts As an extreme parameter-efficient fine-tuning method, Uni-LoRA doesn’t possess any significant limitations other than it is currently evaluated only in English, text-only NLP benchmarks. As for its broader impacts, Uni-LoRA enables high-quality adaptation of LLMs with minimal computational resources, contributing to the democratization of model customization. We do not foresee any societal risks beyond those generally associated with the use of LLMs.
@@ A.1 Representations of additional LoRA/PEFT methods in the unified framework
VB-LoRA learns a shared vector bank =$\mathcal{B}=\{\alpha_{1},\cdots,\alpha_{h}\}$, with $\alpha_{i}\in\mathbb{R}^{b}$, and the compositional coefficients to generate each sub-vector of $B$s and $A$s. Specifically, each sub-vector is generated from top-$K$ (e.g., $K=2$) compositional coefficients over $\mathcal{B}$. Since the coefficients are sparse, we can encode them efficiently by $K$ indices over $\mathcal{B}$ and $K$ coefficient values.
@@
represents the weight increment as $\Delta W=P_{B}\Lambda_{R}P_{A}$, where $P_{B}\in\mathbb{R}^{m\times r}$, $P_{A}\in\mathbb{R}^{r\times n}$ and $\Lambda_{R}\in\mathbb{R}^{r\times r}$ are defined per LoRA module; while $P_{B}$ and $P_{A}$ are initialized and frozen, $\Lambda_{R}$ is trainable. Specifically, for each LoRA module $\ell=1,\cdots,L$, we column-wise flatten each $\Lambda_{R}^{\ell}$, and concatenate them to construct a trainable parameter vector:
@@ A.2 Hyperparameters and Computing Resources
The hyperparameters used for the natural language understanding, mathematical reasoning, and instruction tuning are provided in Tables 78 and 9. All experiments were conducted on a server equipped with 8 NVIDIA A100 80GB GPUs.
@@
The proposed projection in Uni-LoRA is straightforward to implement in modern deep learning frameworks such as PyTorch, enabling full utilization of GPU acceleration. Moreover, since the projection matrix in Uni-LoRA is frozen and does not require training – as opposed to Tied-LoRA and VB-LoRA – it incurs significantly lower computational overhead. As shown in Tables 8 and 9, Uni-LoRA achieves substantially shorter training time compared to VB-LoRA, and is on par with the original LoRA.
@@
Uni-LoRA significantly reduces memory consumption by minimizing the number of trainable parameters. During LoRA fine-tuning, the forward computation is performed as $z=Ax$, $H=Bz$, without the need of explicitly instantiating $\Delta W$. This memory-efficient strategy is seamlessly supported in Uni-LoRA and has been implemented in our codebase. As shown in Tables 8 and 9, Uni-LoRA consistently consumes less memory than both LoRA and VB-LoRA across all model configurations evaluated.
@@
Vector bank $b$ 256 256
Vector bank $h$ 2048 2048
@@
Vector bank $b$ 256 256
Vector bank $h$ 2048 4096
@@
Table 10: All results reported were generated using a consistent version of the GPT-4 API. Score refers to the evaluation on MT-Bench with single-turn dialogues.
@@
A.3 Comparison with LoRA of the Same Rank
In the instruction tuning experiments, we follow the same setup as VeRA [4] and VB-LoRA [6], comparing our method against LoRA with a rank of 64, which is a commonly used setting in prior works. To ensure a fair comparison with our Uni-LoRA, which uses a low rank of 4, we additionally include a set of instruction tuning experiments where LoRA is also configured with rank 4. These experiments follow the same setup as in Section 4.3, with the only difference being the LoRA rank.

### v3

@@ Abstract
Low-Rank Adaptation (LoRA) has become the de facto parameter-efficient fine-tuning (PEFT) method for large language models (LLMs) by constraining weight updates to low-rank matrices. Recent works such as Tied-LoRA, VeRA, and VB-LoRA push efficiency further by introducing additional constraints to reduce the trainable parameter space. In this paper, we show that the parameter space reduction strategies employed by these LoRA variants can be formulated within a unified framework, Uni-LoRA, where the LoRA parameter space, flattened as a high-dimensional vector space $\mathbb{R}^{D}$, can be reconstructed through a projection from a subspace $\mathbb{R}^{d}$, with $d\ll D$. We demonstrate that the fundamental difference among various LoRA methods lies in the choice of the projection matrix, $P\in\mathbb{R}^{D\times d}$. Most existing LoRA variants rely on layer-wise or structure-specific projections that limit cross-layer parameter sharing, thereby compromising parameter efficiency. In light of this, we introduce an efficient and theoretically grounded projection matrix that is isometric, enabling global parameter sharing and reducing computation overhead. Furthermore, under the unified view of Uni-LoRA, this design requires only a single trainable vector to reconstruct LoRA parameters for the entire LLM – making Uni-LoRA both a unified framework and a “one-vector-only” solution. Extensive experiments on GLUE, mathematical reasoning, and instruction tuning benchmarks demonstrate that Uni-LoRA achieves state-of-the-art parameter efficiency while outperforming or matching prior approaches in predictive performance. Our code is available at https://github.com/KaiyangLi1992/Uni-LoRA.
@@ 1 Introduction
To address the aforementioned limitations, we propose Uni-LoRA, a unified framework of LoRA that treats the LoRA parameter space as a high-dimensional vector space and performs a global projection into a shared low-dimensional subspace. We further introduce an efficient and theoretically grounded projection matrix $P\in\mathbb{R}^{D\times d}$, in which each row is a one-hot vector with the index of "1" sampled uniformly from $d$ slots, followed by a column-wise normalization. As discussed in Sec. 3.3, this construction ensures that $P$ possesses three desirable properties for adaptation: globality, uniformity/load-balancing, and isometry. Conceptually, this corresponds to randomly partitioning all the $D$ parameters of LoRA into $d$ groups and enforce the parameters within each group to share the same value during the training process. We prove that the resulting projection matrix is isometric (or distance-preserving). Empirically, when fine-tuning the GEMMA-7B model, our Uni-LoRA achieves performance comparable to LoRA while training only 0.52M parameters – only 0.0061% of the base model size and 0.26% of the LoRA parameter size. Across multiple benchmarks, our method achieves extreme parameter efficiency, while outperforming or matching the state-of-the-art LoRA variants. Moreover, we show that our projection matrix attains the performance of the Fastfood projection (10.5555/3042817.3042964, ). While Fastfood is a widely used structured projection method with a time complexity of $\mathcal{O}(D\log d)$, our method achieves a significantly lower time complexity of $\mathcal{O}(D)$. Our contributions are summarized as follows:
@@ 3.1 Preliminaries: LoRA and its Parameter-efficient Variants
Tied-LoRA renduchintala-etal-2024-tied and VeRA (kopiczko2024vera, ) augment the LoRA architecture with two extra diagonal matrices. Both methods constrain the weight increment as $\Delta W=\Lambda_{b}P_{B}\Lambda_{d}P_{A}$, where $P_{B}\in\mathbb{R}^{m\times r}$ and $P_{A}\in\mathbb{R}^{r\times n}$ are shared/tied across all LoRA modules, and $\Lambda_{b}\in\mathbb{R}^{m\times m}$ and $\Lambda_{d}\in\mathbb{R}^{r\times r}$ (per LoRA module) are trainable diagonal matrices that can selectively enable or disable columns and rows of $P_{B}$ and $P_{A}$ by scaling, allowing effective, fine-grained adaptation with a minimal number of trainable parameters. The main difference between Tied-LoRA and VeRA is that while $P_{B}$ and $P_{A}$ are trainable in Tied-LoRA, they are randomly initialized and frozen in VeRA, leading to a higher parameter efficiency.
@@ 3.3 Analysis of the Projection Matrix
Let $P\in\mathbb{R}^{D\times d}$ be a projection matrix where each row selects exactly one entry uniformly at random to be nonzero, and sets all other entries to zero. Let $n_{j}$ denote the number of nonzero entries in column $j$, and $n_{j}>0$  [^1] To ensure the condition $n_{j}>0$ is always held, we can re-sample $P$ if one column happens to be all zeros.. For column-wise normalization, each nonzero entry in column $j$ is set to $1/\sqrt{n_{j}}$. As a projection matrix, $P$ is isometric. Formally, $\|P(x-y)\|=\|x-y\|,\forall x,y\in\mathbb{R}^{d}$.
@@
Given the construction of $P$ in the theorem statement, we begin by showing that $P^{\top}P=I_{d}$. Consider the $(j,k)$-th entry of $^{\top}P: [PP$$^{\top}P]_{j,k}=\sum_{i=1}^{D}P_{i,j}P_{i,k}$. Case 1: $j\neq k$. Since each row contains only one nonzero entry, there exists no row $i$ such that both $P_{i,j}$ and $P_{i,k}$ are nonzero. Hence, every term in the summation is zero, and we have $P^{\top}P]_{j,k}=0$. Case 2: $j=k$. Column $j$ contains $n_{j}$ nonzero entries, each of value $1/\sqrt{n_{j}}$. Thus, $P^{\top}P]_{j,j}=\sum_{i=1}^{D}P_{i,j}^{2}=n_{j}\cdot\left({1}/{\sqrt{n_{j}}}\right)^{2}=1$. From the analysis above, we have $P^{\top}P]_{j,k}=1$, if $j=k$ and 0 otherwise, thus $P^{\top}P=I_{d}$. Therefore,
@@ 4 Experiments
In this section, we evaluate our method through a series of experiments. We begin by comparing Uni-LoRA to the state-of-the-art PEFT methods: LoRA, Tied-LoRA, VeRA, LoRA-XS, and FourierFT on the GLUE benchmark. Next, we extend our analysis to mathematical reasoning tasks on Mistral and Gemma models, as well as instruction tuning tasks on Llama2. All our experiments are conducted on a server equipped with 8 NVIDIA A100 80GB GPUs. For reproducibility, we provide detailed hyperparameters and specifications of computing resources for each experiment in Appendix A.2.
@@ 4.2 Mathematical Reasoning
To evaluate mathematical reasoning capabilities, we fine-tune the Mistral-7B-v0.1 (jiang2023mistral7b, ) and Gemma-7B (gemmateam2024gemmaopenmodelsbased, ) models on the MetaMathQA (yu2023metamath, ) dataset and test them on GSM8K (cobbe2021gsm8k, ) and MATH (hendrycksmath2021, ). We compare our results with the state-of-the-art methods, following their experimental configurations. Table 4 shows that Uni-LoRA achieves strong and consistent performance across both Mistral-7B and Gemma-7B backbones. On the MATH benchmark, it reaches 18.18 with Mistral-7B and 28.94 with Gemma-7B, outperforming LoRA-XS, VB-LoRA and FourierFT. It also yields higher accuracy than LoRA-XS on both GSM8K and MATH, while using less than 70% of its trainable parameters. Compared to VB-LoRA, Uni-LoRA achieves similar or better performance with substantially fewer parameters. It also demonstrates comparable accuracy to VeRA on mathematical reasoning tasks, while requiring only half as many trainable parameters on Mistral-7B and less than one-third on Gemma-7B. Furthermore, Uni-LoRA consistently outperforms FourierFT on both tasks, while being more parameter-efficient.
Table 3: Results with the Mistral-7B and Gemma-7B models on the GSM8K and MATH Benchmarks. The best results in each group are shown in bold. Baselines’ results are either reproduced using their reported configurations or directly sourced from their original papers.
@@ 4.3 Instruction Tuning
4.4 Experiments of Computer Vision Tasks
To further assess the effectiveness of Uni-LoRA on more complex vision tasks, we follow the experimental setup in FourierFT (DBLP:conf/icml/GaoWCLWC024, ) and conduct additional evaluations. Specifically, we apply Uni-LoRA with d=74,000d=74{,}000 on ViT-Base and d=144,000d=144{,}000 on ViT-Large, using a fixed rank $r=4$. A grid search is conducted over the head learning rate in $\{1\!\times\!{10^{-3}},\ 2\!\times\!10^{-3},\ 5\!\times\!10^{-3},\ 1\!\times\!10^{-2}\}$ and the trainable vector learning rate in $\{2\!\times\!10^{-3},\ 5\!\times\!10^{-3},\ 1\!\times\!10^{-2},\ 2\!\times\!10^{-2},\ 5\!\times\!10^{-2}\}$. The number of training epochs is fixed at 20. All experiments are repeated five times, with the mean and standard deviation across the runs reported in Table LABEL:tab:uni_lora_cv.
Table 5: Comparison of Uni-LoRA and baseline methods on eight computer vision datasets using ViT-Base and ViT-Large backbones. All baseline results, including Full Fine-tuning ($F$) and Linear Probing (LP), are taken from the original FourierFT paper DBLP:conf/icml/GaoWCLWC024 . In LP, only the classification head is fine-tuned.
Model Method # Trainable OxfordPets StanfordCars CIFAR10 DTD EuroSAT FGVC RESISC45 CIFAR100 Avg.
Para.
[…]
Experiments show that Uni-LoRA consistently outperforms FourierFT across computer vision benchmarks. Remarkably, even when FourierFT uses three times as many trainable parameters, Uni-LoRA still achieves better performance. Furthermore, when the number of trainable parameters in Uni-LoRA is less than one-thousandth of that in full fine-tuning, the performance gap between Uni-LoRA and full fine-tuning remains within 2.5%. These results highlight the strong efficiency and generalizability of Uni-LoRA in complex computer vision tasks.
@@ 4.5 Ablation Study
Fastfood (10.5555/3042817.3042964, ) is a classical projection method known for its low computational and memory cost and distance-preserving properties. To evaluate the effectiveness of our projection method, we compare it against Uni-LoRA (Fastfood) on four GLUE tasks in terms of both predictive performance and runtime efficiency. Table 7 shows that Uni-LoRA with the uniform random projection consistently outperforms Uni-LoRA (Fastfood) in terms of predictive performance across four GLUE tasks, while being significantly faster. For example, on MRPC, it achieves a slightly higher score (91.3 vs. 90.7) with a 65% reduction in training time (9 vs. 26 mins). This efficiency stems from the linear time complexity $\mathcal{O}(D)$ of our projection, compared to Fastfood’s $\mathcal{O}(D\log d)$.
@@
To investigate the impact of global vs. layer-wise (local) projections to our method, we construct a controlled variant as follows. On the RoBERTa-large model, we project the parameters of each layer into its own trainable subspace, with each local projection matrix constructed in the same way as that of Uni-LoRA’s. To ensure a fair comparison, we set the per-layer subspace dimensionality such that the total number of trainable space dimensionality matches that of Uni-LoRA’s. Table 7 shows that local projection performs worse than global projection on MRPC, SST-2, and QNLI, while both methods achieve comparable performance on CoLA.
@@
To investigate the impact of uniform vs. non-uniform projections to our method, we construct a controlled variant in which two separate projection matrices are used: one projects all LoRA $A$ matrices into the first two-thirds of the subspace dimensions, and the other projects all $B$ matrices into the remaining one-third. Both projection matrices are constructed using the same method as that of Uni-LoRA’s. Table 7 shows that the uniform projection consistently outperforms the non-uniform projection across all the tasks.
@@ 5 Conclusion
Limitations and broader impacts As an extreme parameter-efficient fine-tuning method, Uni-LoRA doesn’t possess any significant limitations other than it is currently only evaluated on small to medium scale NLP/CV benchmarks due to limited computing resources. As for its broader impacts, Uni-LoRA enables high-quality adaptation of LLMs with minimal computational resources, contributing to the democratization of model customization. We do not foresee any societal risks beyond those generally associated with the use of LLMs.
Acknowledgments
We would like to thank the anonymous reviewers for their excellent comments and suggestions, which greatly helped improve the quality of this paper. This work was supported in part by the National Science Foundation under Grant Nos. 2343619, 2416872, 2244219, 2315596, and 2146497.
@@ A.1 Representations of additional LoRA/PEFT methods in the unified framework
VB-LoRA learns a shared vector bank $\mathcal{B}=\{\alpha_{1},\cdots,\alpha_{h}\}$, with $\alpha_{i}\in\mathbb{R}^{b}$, and the compositional coefficients to generate each sub-vector of $B$s and $A$s. Specifically, each sub-vector is generated from top-$K$ (e.g., $K=2$) compositional coefficients over $\mathcal{B}$. Since the coefficients are sparse, we can encode them efficiently by $K$ indices over $\mathcal{B}$ and $K$ coefficient values.
@@
represents the weight increment as $\Delta W=P_{B}\Lambda_{R}P_{A}$, where $P_{B}\in\mathbb{R}^{m\times r}$ and $P_{A}\in\mathbb{R}^{r\times n}$ are derived from the singular value decomposition (SVD) of the pretrained weight matrix $W$ and are kept frozen during fine-tuning, while $\Lambda_{R}\in\mathbb{R}^{r\times r}$ is the only trainable component. Specifically, for each LoRA module $\ell=1,\cdots,L$, we column-wise flatten each $\Lambda_{R}^{\ell}$, and concatenate them to construct a trainable parameter vector:
@@ A.2 Hyperparameters and Computing Resources
The hyperparameters used for the natural language understanding, mathematical reasoning, and instruction tuning are provided in Tables 89,  10 and 11 . All experiments were conducted on a server equipped with 8 NVIDIA A100 80GB GPUs.
@@
The proposed projection in Uni-LoRA is straightforward to implement in modern deep learning frameworks such as PyTorch, enabling full utilization of GPU acceleration. Moreover, since the projection matrix in Uni-LoRA is frozen and does not require training – as opposed to Tied-LoRA and VB-LoRA – it incurs significantly lower computational overhead. As shown in Tables 9 and 10, Uni-LoRA achieves substantially shorter training time compared to VB-LoRA, and is on par with the original LoRA.
@@
Uni-LoRA significantly reduces memory consumption by minimizing the number of trainable parameters. During LoRA fine-tuning, the forward computation is performed as $z=Ax$, $H=Bz$, without the need of explicitly instantiating $\Delta W$. This memory-efficient strategy is seamlessly supported in Uni-LoRA and has been implemented in our codebase. As shown in Tables 9 and 10, Uni-LoRA consistently consumes less memory than VB-LoRA across all model configurations evaluated.
@@
Vector bank $b$ - - 256 256
Vector bank $h$ - - 2048 2048
@@
Vector bank $b$ - - - - 256 256
Vector bank $h$ - - - - 2048 4096
@@
Table 11: Hyperparameters and computing resources used in the vision tasks. h: hour, m: minute.
Model Hyperparameter OxfordPets StanfordCars CIFAR10 DTD EuroSAT FGVC RESISC45 CIFAR100
Optimizer AdamW
Weight Decay 0.01
LR Schedule Linear
Init. of $\theta_{d}$ $\mathcal{U}(-0.02,0.02)$
Epoch 20
Base # GPUs 1
Learning Rate (Head) 5E-3 5E-2 5E-2 5E-2 5E-2 5E-2 5E-2 1E-2
Learning Rate ($\theta_{d}$) 5E-3 1E-2 1E-2 1E-2 1E-2 1E-2 1E-2 1E-2
$|\theta_{d}|$ 72,000
Rank 4
Batch Size Per GPU 128
Training Time 6m 23m 49m 13m 40m 53m 30m 1.6h
GPU Memory 12,132 MiB
Large # GPUs 1
Learning Rate (Head) 2E-2 1E-2 5E-3 1E-2 2E-2 1E-2 1E-2 5E-3
Learning Rate ($\theta_{d}$) 1E-2 1E-2 1E-2 1E-2 1E-2 1E-2 1E-2 1E-2
$|\theta_{d}|$ 144,000
Rank 4
Batch Size Per GPU 32
Training Time 11m 33m 1.6h 13m 40m 53m 50m 1.6h
GPU Memory 30,058 MiB
Table 12: All results reported are generated using a consistent version of the GPT-4 API. Score refers to the evaluation on MT-Bench with single-turn dialogues.
@@
A.3 Impact of Hyperparameters on Model Performance
To investigate how the number of trainable parameters $d$ affects model performance, we conducted ablation studies by fine-tuning RoBERTa-Large on the SST-2 task (from GLUE) and Gemma-7B on the mathematical reasoning benchmarks, where $d$ is varied while keeping all other settings at their default values. The results are summarized in Figure 3.
(a) Accuracy on SST-2 as $d$ increases.
(b) Accuracies on GSM8K and MATH as $d$ increases.
Figure 3: Evolution of Uni-LoRA accuracies on different benchmarks as the number of trainable parameters $d$ increases.
Our experiments show that performance improves rapidly with increasing $d$ when $d$ is small, and then plateaus as $d$ grows larger.
For a fair comparison, we follow VB-LoRA and set the LoRA rank $r$ = 4 in the paper. To evaluate the influence of rank $r$, here we conduct additional analysis by varying $r$ for RoBERTa-Large on the SST-2 task (from GLUE) and Gemma-7B on the mathematical reasoning benchmarks. The results reported in Figure 4 show that Uni-LoRA maintains stable performance across a wide range of ranks, with $r$ = 4 achieving the best balance between accuracy and efficiency.
(a) Accuracy on SST-2 with different $r$.
(b) Accuracies on GSM8K and MATH with different $r$.
Figure 4: Performance comparison of Uni-LoRA across different rank $r$.
A.4 Comparison with LoRA of the Same Rank
In the instruction tuning experiments, we follow the same setup in VeRA [4] and VB-LoRA [6], comparing our method against LoRA with a rank of 64, which is a commonly used setting in prior works. To ensure a fair comparison with our Uni-LoRA, which uses a low rank of 4, we additionally include a set of instruction tuning experiments where LoRA is also configured with rank 4. These experiments follow the same setup as in Section 4.3, with the only difference being the LoRA rank.